# Gas-induced perturbations on the gravitational wave in-spiral of live post-Newtonian LISA massive black hole binaries

### v1

@@ Abstract
We investigate the effect of dynamically coupling gas torques with gravitational wave (GW) emission during the orbital evolution of an equal-mass massive black hole binary (MBHB). We perform hydrodynamical simulations of eccentric MBHBs with total mass $M=10^{6}~{}{\rm M}_{\sun}$ embedded in a prograde locally isothermal circumbinary disk (CBD). We evolve the binary from $53$ to $30$ Schwarzschild radii separations using up to 2.5 post-Newtonian (PN) corrections to the binary dynamics, which allow us to follow the GW-driven in-spiral. For the first time, we report the measurement of gas torques onto a live binary a few years before the merger, with and without concurrent GW radiation. We also identify and measure a novel GW-gas coupling term in the in-spiral rate that makes gas effects an order of magnitude stronger than the gas-only contribution. We show that the evolution rate ($\dot{a}$) of the MBHB can be neatly expressed as the sum of the GW rate ($\dot{a}_{\rm GW}$), the pure gas-driven rate ($\dot{a}_{\rm gas}$), and their cross-term $\propto\dot{a}_{\rm GW}\dot{a}_{\rm gas}$. The source-frame gas-induced dephasing in the GW waveform is equivalent to losing $\sim 0.5$ GW cycles over the expected $\sim 1700$ cycles in a vacuum, which LISA should detect at redshift $z=1$. We also propose a phenomenological model that captures the essence of simulations and can be used to perform Bayesian inference. Our results show how GWs alone can be used to probe the astrophysical properties of CBDs and have important implications on multi-messenger strategies aimed at studying the environments of MBHBs.
@@ 1 Introduction
The recent adoption of LISA (Amaro-Seoane et al., 2017; Colpi et al., 2024) and the development of TianQin (Li et al., 2024) and Taiji (Gong et al., 2021) will provide a powerful opportunity to detect gravitational waves (GWs) from coalescing near-equal mass massive black hole binaries (MBHBs) with masses $\sim 10^{4}$-$10^{7}~{}{\rm M}_{\sun}$. LISA can potentially detect MBHBs up to redshifts $z\lesssim 20$ and with high (e.g. $\lesssim 10^{3}$) signal-to-noise ratios (SNRs; Amaro-Seoane et al., 2017). MBHBs are a by-product of galaxy mergers (Begelman et al., 1980). When two galaxies merge, the massive black holes (MBHs) hosted in their centre are expected to reach the centre of the remnant galaxy owing to the dynamical friction mechanism and form a bound binary at a pc scale. This binary can further proceed towards merger through the interaction with surrounding stars and gas until GWs are strong enough to take over and drive the binary to coalescence (Amaro-Seoane et al., 2023).
@@
While interactions with stars in a tri-axial potential (e.g. Quinlan, 1996; Preto et al., 2011; Khan et al., 2011) as well as a third close-by MBH (Blaes et al., 2002; Hoffman & Loeb, 2007; Bonetti et al., 2019) can lead to a MBHB merger, they are relatively slow and rare mechanisms, respectively. In particular, sinking timescales of MBHBs due to three body encounters with stars can exceed a Gyr in the low density environments of stellar-disk dominated galaxies (Khan et al., 2018), the typical hosts of MBHs below $10^{6}~{}{\rm M}_{\sun}$, that fall in the mass range accessible by LISA. On the other hand, when the host galaxies are gas-rich, and have circumnuclear gas disks, then MBHBs can sink efficiently below pc separations (Mayer, 2013; Souza Lima et al., 2020). At even smaller separations, below $\sim 0.1$ pc, the MBHB would carve a cavity, which results in the formation of a co-planar geometrically thin (Shakura & Sunyaev, 1973) circumbinary disk (CBD; Escala et al. 2004; Cuadra et al. 2009; D’Orazio et al. 2016), allowing the binary to coalesce in less than $100$ Myr (Haiman et al., 2009). Over the past few years, the interaction between a binary and its circumbinary disk has been studied extensively for various system parameters and thermodynamics assumptions using several different numerical hydrodynamical (HD) simulations (see e.g. Duffell et al., 2024). The general consensus is that circular nearly equal-mass binaries do undergo out-spiral in relatively thick CBDs while their in-spiral is aided by relatively thin (i.e. aspect ratio $H/R\lesssim 0.03$) CBDs (Tiede et al., 2020; Franchini et al., 2021, 2022).
@@
Since MBHBs observable by LISA are likely to reside in gaseous environments (see, e.g. Mangiagli et al., 2022), it is important to study the effect of gas on the orbital evolution of the MBHB when it enters the LISA band. The first attempt in this direction was performed by several groups (Garg et al., 2022; Tiede et al., 2024; Zwick et al., 2024; Garg et al., 2024a) who measured the gas-induced dephasing in the LISA band by simply linearly adding the gas-driven evolution rate, computed in post-processing from 2D HD fixed binary orbit simulations, to the GW in-spiral rate. However, the scales considered in those numerical works are close to sub-pc, where GWs are still too weak to drive significant binary evolution. Furthermore, by adding the two contribution linearly, gas-induced dephasing studies might have ignored possible coupling between gas torques and GW-driven evolution, due to the lack of HD simulations where the two effects are naturally coupled together and the binary evolves under both processes at the same time.
@@
Recently, Franchini et al. (2024) simulated an eccentric, live (Franchini et al., 2023), equal-mass $10^{6}~{}{\rm M}_{\sun}$ MBHB embedded in a prograde $100{\rm M}_{\sun}$ CBD by dynamically modeling the binary in-spiral with post-Newtonian (PN) corrections up to $2.5$ order. They evolved the system for the final years of in-spiral, including the merger and post merger phase, to quantify possible electromagnetic (EM) counterparts. In this work, we use the same setup to simulate the same binary but now embedded in a lighter $5~{}{\rm M}_{\sun}$ disk to properly investigate how gas perturbs the binary evolution rate. We then quantify for the first time the effect of gas-induced perturbations on waveforms using a live binary whose dynamics is computed using PN corrections, thus including the interplay between energy and angular momentum change caused by both GW radiation and gas torques. With this simulation setup, the in-spiral is concurrently determined by GWs and gas, as opposed to co-adding the two effects in post-processing as previously done in the literature, allowing us to robustly quantify the gas-induced dephasing in the GW waveform and its detectability by LISA.
@@ 2 Numerical setup
Following the approach used in Franchini et al. (2024), we model the binary using two equal-mass sink particles (Bate et al., 1995) that represent two Schwarzschild MBHs with total mass $M=10^{6}~{}{\rm M}_{\sun}$ , setting the respective sink radius to the innermost stable circular orbit (ISCO). We set the MBHB initial semi-major axis (SMA) to $a=53.2$ Schwarzschild radii ($r_{s}$) and eccentricity to $e=0.29$. Note that we start from the initial condition of the thin (i.e. aspect ratio $H/R=0.03$) disk simulation in Franchini et al. (2024), which originated from a circular equal-mass binary evolved for $1000$ binary orbits by Franchini et al. (2022). During the first $1000$ orbits the eccentricity of the simulated live binary increases to $e=0.29$ as a result of the interaction with the CBD. We here assume the disk to have a much smaller mass $M_{\rm d}=5{\rm M}_{\sun}$.
@@
We follow the evolution of the binary driven by both gas torques and PN corrections up to 2.5PN order using the code gizmo (Hopkins, 2015) until the binary reaches $=30\,r_{\rm s}$ in separation in our GW+gas simulation, i.e. for $\sim 1700$ GW cycles. The implementation of the PN corrections to the binary dynamics follows the equations in Blanchet (2014). We include both conservative $1$PN and $2$PN terms, and radiative $2.5$PN terms. The latter term generates the GW emission and leads to the decrease in binary SMA ($\dot{a}_{\rm GW}$) and eccentricity only due to GWs. In order to integrate the 2.5PN equations, we implemented an intermediate predictor step to update the particle velocities at the end of the time step, accounting for the PN corrections, and re-enforcing the numerical stability of the integration algorithm. Our approach is similar to the one outlined in Sect. 6.2 of Liptai & Price (2019), except that we use a predictor-corrector approach instead of implementing the implicit kick-drift-kick one. In the initial setup the disk extended from $2a$ to $10a$. However during the first 1000 binary orbits, the cavity becomes eccentric and the inner edge increases to $\sim 3.5a$.
We also run two other simulations with the same setup. First, without the dissipative $2.5$PN term to consider the binary evolution only due to the presence of the CBD. We run it for $1000$ orbits to see any appreciable changes in the orbital quantities. This allows us to infer the SMA evolution rate only due to the gaseous disk ($\dot{a}_{\rm gas}$) and therefore to disentangle its effect from the putative GW-gas cross term considered later. We then also run a simulation without the CBD in order to self-consistently obtain the evolution due to the PN terms only. This allows us to mitigate the numerical error in the integration when we compute the difference between the simulation with concurrent evolution under GWs and gas, and the one with PN terms only.
@@ 2.1 Post-processing analysis
The CBD typically affects the binary evolution by exerting both a gravitational torque ($T_{\rm grav}$) and an accretion torque ($T_{\rm acc}$). The first, just due to gravity, is essentially driven by any asymmetry in the gaseous flow while the latter is instead induced by the accretion of gas particles onto either MBHs. In particular, the accretion of gas alters not only the mass but also the angular momentum of the binary (see Franchini et al. 2021 for detailed calculations).
We can then express the overall gas effect in terms of a single simulation-calibrated parameter (Garg et al., 2022; Duffell et al., 2024)
@@
where $\dot{M}\equiv 0.02~{}{\rm f}_{\rm Edd}(M/10^{6}{\rm M}_{\sun}){({\rm M}_{\sun}/\rm yr)}$ is the accretion rate onto the binary for our assumed $\epsilon=0.1$ radiative efficiency, $\Omega\equiv\sqrt{GM/a^{3}}$ is the binary orbital angular frequency, and $\dot{M}a^{2}\Omega$ is the viscous torque (Lin & Papaloizou, 1986). $\xi$ depends sensitively on the binary and disk parameters, in particular on the binary mass ratio $q$, disk shape and temperature, and may also depends upon the assumed equation of state. Previous sub-pc fixed binary orbit 2D Newtonian simulation predicted $|\xi|\sim 3$ (Dittmann & Ryan, 2022) for $H/R\sim 0.03$.
We can therefore write down the SMA evolution rate due to gas torques for a quasi-circular binary as (Garg et al., 2024a)
$\dot{a}_{\rm gas}=8\xi\frac{\dot{M}_{z}}{M_{z}}a=\xi{\rm f}_{\rm Edd}\frac{4}{25{\rm Myr}}a,$ (2)
where we infer both $\xi$ and ${\rm f}_{\rm Edd}$ from gas-only simulation to properly isolate any GW-gas coupling in the total SMA rate in the GW+gas simulation. Moreover, since the evolution only due to gas torques changes the SMA by a small quantity even after $1000$ orbits, we extrapolate the results to smaller separations, i.e. until $30r_{s}$.
When the MBHB is dynamically evolving under both GWs and gas, there can be a coupled evolution that manifests in the form of a cross term $\dot{a}_{\rm GWgas}$, whose exact functional form is unknown. Therefore, the total SMA evolution rate can be written down as
$\dot{a}_{\rm GW+gas}=\dot{a}_{\rm GW}+\dot{a}_{\rm gas}+\dot{a}_{\rm GWgas},$ (3)
where both $\dot{a}_{\rm gas}$ and $\dot{a}_{\rm GWgas}$ can be thought of as perturbative corrections to $\dot{a}_{\rm GW}$ in the final years of in-spiral.
Assuming the stationary phase approximation (Cutler & Flanagan, 1994), where gas affects the phase but not the amplitude of the GW, we compute the dephasing in the time-domain GW waveform, caused by the presence of the CBD as
$\delta\phi_{\rm gas}=\Phi(\dot{a}_{\rm GW+gas})-\Phi(\dot{a}_{\rm GW}),$ (4)
where negative (positive) dephasing imply that gas speed-up (slow-down) the binary in-spiral. Here $\Phi(\dot{a})$ is the total accumulated GW phase for a given evolution rate $\dot{a}$ between an initial ($a_{i}$) and a final ($a_{f}$) SMAs, and it can be expressed in radians as
$\Phi(\dot{a})=2\pi\int_{a_{\rm i}}^{a_{\rm f}}{\rm d}a\frac{f}{\dot{a}},$ (5)
where $f\equiv\Omega/\pi$ is the GW frequency, that is twice the orbital frequency ($\Omega/2\pi$), under the assumption of quasi-circular binaries.
@@ 3 Results
We show disk morphologies and associated column densities at three SMAs – $a=50r_{s}$, $a=40r_{s}$, and $a=30r_{s}$ – in Fig. 1 for the GW+gas simulation.111disk morphology for gas-only simulation is similar to the left panel of Fig. 1. The gas morphology is similar, as expected, to the simulations presented in Franchini et al. (2024). We can clearly see the over-density at the cavity edge, i.e. the lump (Shi et al., 2012), precessing around the binary.
@@
We measure the accretion rate in both our simulations and find a mean Eddington ratio of ${\rm f}_{\rm Edd}\approx 0.92$ for GW+gas simulation and ${\rm f}_{\rm Edd}\approx 1.20$ for the gas-only case. The slightly higher accretion rate in the gas-only case is a natural consequence of the fact that the binary does not start decoupling from the disk because of the lack of GW radiation, therefore the gas can flow inside the cavity and keep feeding the MBHB more effectively.
@@
Figure 1: Column density ($\Sigma$) plots at three SMAs: $50r_{s}$ (left panel), $40r_{s}$ (middle panel), and $30r_{s}$ (right panel) for the binary evolution under both GW and gas. Here $\Sigma$ varies between $\sim 10^{3.5}$-$10^{6}$ g/cm2. While both the binary (green dots) and the cavity shrinks with time, cavity is becoming more eccentric also.
We then compute both gravitational and accretion torques directly from the simulations. We find that the magnitude of the accretion torque is $T_{\rm acc}\sim 10^{-2}T_{\rm grav}$. We show the evolution of the parameter $\xi$ in Fig. 2 for GW+gas simulations as a function of SMA together with the average values of $\xi$ for both GW+gas ($\bar{\xi}_{\rm GW+gas}\approx-24$) and gas-only ($\bar{\xi}_{\rm gas}\approx-36$) cases. Unsurprisingly, the mean gas effect is stronger when the binary is only evolving due to the gas torque. The parameter $\xi$ is indeed decreasing in magnitude as the binary is in-spiraling further, since the gas effects are becoming weaker. Since the value of $\xi$ does not change significantly with time in the gas-only simulation, the approximation we make in the parametrization of Eq. (1) is reasonable. It is worth noticing that for the computed gas-induced dephasing $\delta\phi_{\rm gas}$ (see Section 3.1), most of the contribution is expected to originate from the early in-spiral phase, where the mean value of $\xi$ is higher.
We also note that the high-frequency or sub-orbital fluctuations in Fig. 2 around the mean value may be measurable on their own, as suggested by analytical studies (Zwick et al., 2021, 2024). However, we focus on the secular trend in this work.

### v3

@@ Abstract
We investigate the effect of dynamically coupling gas torques with gravitational wave (GW) emission during the orbital evolution of an equal-mass massive black hole binary (MBHB). We perform hydrodynamical simulations of eccentric MBHBs with total mass $M=10^{6}~{\rm M}_{\sun}$ embedded in a prograde locally isothermal circumbinary disk (CBD). We evolve the binary from $55$ to $49$ Schwarzschild radii separations using up to 2.5 post-Newtonian (PN) corrections to the binary dynamics, which allow us to follow the GW-driven in-spiral. For the first time, we report the measurement of gas torques onto a live binary a few years before the merger, with and without concurrent GW radiation. We also report the gas-induced orbital dephasing $\delta\phi_{\rm orb}\sim-0.007$ rad over $278$ orbital cycles that is likely driven mainly by disk-induced precession and LISA should be able to detect it at redshift $z=1$. Our results show how GWs alone can be used to probe the astrophysical properties of CBDs and have important implications for multi-messenger strategies aimed at studying the environments of MBHBs.
@@ 1 Introduction
The recent adoption of LISA (Amaro-Seoane et al., 2017; Colpi et al., 2024) and the development of TianQin (Li et al., 2025) and Taiji (Gong et al., 2021) will provide a powerful opportunity to detect gravitational waves (GWs) from coalescing near-equal mass massive black hole binaries (MBHBs) with masses $\sim 10^{4}$-$10^{7}~{\rm M}_{\sun}$. LISA can potentially detect MBHBs up to redshifts $z\lesssim 20$ and with high (e.g. $\lesssim 10^{3}$) signal-to-noise ratios (SNRs; Amaro-Seoane et al., 2017). MBHBs are a by-product of galaxy mergers (Begelman et al., 1980). When two galaxies merge, the massive black holes (MBHs) hosted in their centre are expected to reach the centre of the remnant galaxy owing to the dynamical friction mechanism and form a bound binary at pc scales. This binary can further proceed towards merger through the interaction with surrounding stars and gas until GWs are strong enough to take over and drive the binary to coalescence (Amaro-Seoane et al., 2023).
@@
While interactions with stars in a tri-axial potential (e.g. Quinlan, 1996; Preto et al., 2011; Khan et al., 2011) as well as a third close-by MBH (Blaes et al., 2002; Hoffman & Loeb, 2007; Bonetti et al., 2019) can lead to a MBHB merger, they are relatively slow and rare mechanisms, respectively. In particular, sinking timescales of MBHBs due to three body encounters with stars can exceed a Gyr in the low density environments of stellar-disk dominated galaxies (Khan et al., 2018), the typical hosts of MBHs below $10^{6}~{\rm M}_{\sun}$, that fall in the mass range accessible by LISA. On the other hand, when the host galaxies are gas-rich, and have circumnuclear gas disks, then MBHBs can sink efficiently below pc separations (Mayer, 2013; Souza Lima et al., 2020). At separations below $\sim 0.1$ pc, the gas dragged by the two MBHs is expected to settle in a common circumbinary disk (CBD; Escala et al. 2004; Cuadra et al. 2009; D’Orazio et al. 2016), whose torques remove angular momentum from the binary allowing it to coalesce in less than $100$ Myr (Haiman et al., 2009). The binary potential will open a cavity inside the CBD whose size depends on the disk properties (Artymowicz & Lubow, 1994, 1996). Over the past few years, the interaction between a binary and its circumbinary disk has been studied extensively for various system parameters and thermodynamics assumptions using different numerical hydrodynamical (HD) simulations (see e.g. Duffell et al., 2024). The general consensus is that circular nearly equal-mass binaries do undergo out-spiral in relatively thick CBDs while their in-spiral is aided by relatively thin (i.e. aspect ratio $H/R\lesssim 0.03$) CBDs (Tiede et al., 2020; Franchini et al., 2021, 2022).
@@
Since MBHBs observable by LISA are likely to reside in gaseous environments (see, e.g. Mangiagli et al., 2022), it is important to study the effect of gas on the orbital evolution of the MBHB when it enters the LISA band. The first attempt in this direction was performed by several groups (Garg et al., 2022; Tiede et al., 2024; Zwick et al., 2024; Garg et al., 2024a, c, b; Dittmann et al., 2023) who measured the gas-induced dephasing in the LISA band by simply linearly adding the gas-driven evolution rate, computed in post-processing from 2D HD fixed binary orbit simulations, to the GW in-spiral rate. However, the scales considered in those numerical works are close to sub-pc, where GWs are still too weak to drive significant binary evolution. Furthermore, by adding the two contribution linearly, gas-induced dephasing studies might have ignored possible coupling between gas torques and GW-driven evolution, due to the lack of HD simulations where the two effects are naturally coupled together and the binary evolves under both processes at the same time.
@@
Recently, Franchini et al. (2024) simulated an eccentric, live (Franchini et al., 2023), equal-mass $10^{6}~{\rm M}_{\sun}$ MBHB embedded in a prograde $100~{\rm M}_{\sun}$ CBD by dynamically modeling the binary in-spiral with post-Newtonian (PN) corrections up to $2.5$ order. They evolved the system for the final years of in-spiral, including the merger and post-merger phase, to quantify possible electromagnetic (EM) counterparts. In this work, we use the same setup to simulate the same binary but now embedded in a lighter $5~{\rm M}_{\sun}$ disk to properly investigate how gas perturbs the binary evolution rate. We then quantify for the first time the effect of gas-induced perturbations on waveforms using a live binary whose dynamics is computed using PN corrections, thus including the interplay between energy and angular momentum change caused by both GW radiation and gas torques. With this simulation setup, the in-spiral is concurrently determined by GWs and gas, as opposed to co-adding the two effects in post-processing as previously done in the literature, allowing us to robustly quantify the gas-induced dephasing in the GW waveform and its detectability by LISA.
@@ 2 Numerical setup
[…]
We explore three resolutions for our gas+PN2.5 setup by increasing the number of splitting levels in the hyper-Lagrangian refinement. We quantify the resolution in terms of inter-particle spacing $\Delta x$ at $R=3a$. We label these simulations low-resolution (LR) with $\Delta x[3a]={1.31~r_{s}}$, mid-resolution (MR) with $\Delta x[3a]={1.07~r_{s}}$, and high-resolution (HR) with $\Delta x[3a]={0.66~r_{s}}$. We run at least $100$ orbits for each simulations in order to perform a meaningful resolution study, which we report in Appendix A. We find that the gravitational torque exerted by the disk onto the binary in the MR simulation is already converged and therefore we consider the MR run as our fiducial setup. Unless stated otherwise, the following results have been inferred using our MR run.
@@
We follow the evolution of the binary driven by both gas torques and PN corrections up to 2.5PN order using the code gizmo (Hopkins, 2015) until the binary reaches $48.9\,r_{\rm s}$ in separation in our gas+PN2.5 simulation, i.e. for $278$ initial binary orbits or $\sim 556$ GW cycles. The implementation of the PN corrections to the binary dynamics follows the equations in Blanchet (2014). We include both conservative 11PN and 22PN terms, and radiative 2.52.5PN terms. The latter term generates the GW emission and leads to the decrease in binary SMA ($\dot{a}_{\rm GW}$) and eccentricity only due to GWs. In order to integrate the 2.5PN equations, we implemented an intermediate predictor step to update the particle velocities at the end of the time step, accounting for the PN corrections, and re-enforcing the numerical stability of the integration algorithm. Our approach is similar to the one outlined in Sect. 6.2 of Liptai & Price (2019), except that we use a predictor-corrector approach instead of implementing the implicit kick-drift-kick one (see Franchini et al. (2024) for more details).
In order to measure the effect of the gas contribution in the absence of GW emission, we run a simulation with the CBD, but without the 2.5PN dissipative term in the binary orbital motion. We refer to this simulation with the term “gas+PN2” and we perform it only at mid-resolution, i.e. with $\Delta x[3a]={1.07~r_{s}}$. We run this simulation for the same time as the gas+PN2.5 run to see any appreciable changes in the orbital quantities due to the sole interaction with the gaseous disk. This allows us to infer the gas torques and gas-induced orbital as well as binary precession dephasings without the effects introduced by GWs dissipation. Note that we extrapolate the mean torque results of the gas+PN2 simulation over the same integrated time down to $48.9~r_{\rm s}$ since evolving the binary under the mere effect of gas to such small separations is currently computationally prohibitive. Similarly, we compare the binary evolution between different setups over the same integrated time to compute the dephasings.
We then also run two simulations without the CBD in order to self-consistently obtain the binary evolution driven only by the PN terms. We run one simulation including all the PN corrections, including the dissipative GW term, and another simulation with only the 1+2 PN corrections. We refer to the first and second simulation with the label “PN2.5” and “PN2” respectively. These simulations allow us to isolate the effects of GW emission and of the interaction with the disk, and to mitigate numerical errors in the integration when we compute the difference between the simulation with gas and their non-gaseous counterparts.
@@ 2.1 Post-processing analysis
The CBD affects the binary evolution by exerting both a gravitational torque ($T_{\rm grav}$) and an accretion torque ($T_{\rm acc}$). The first, just due to gravity, is essentially driven by any asymmetry in the gaseous flow while the latter is instead induced by the accretion of gas particles onto either MBHs. In particular, the accretion of gas alters not only the mass but also the angular momentum of the binary (see Franchini et al. 2021 for detailed calculations).
We can then express the overall gas effect in terms of a single dimension-less simulation-calibrated parameter (similar to the accretion eigenvalue mentioned in Duffell et al. 2024)
@@
Here $\Omega\equiv\sqrt{GM/a^{3}}$ is the binary orbital angular frequency, and $\dot{M}a^{2}\Omega$ is simply the normalization commonly used in the literature (see, e.g. Duffell et al., 2024). $\xi$ depends sensitively on the binary and disk parameters, in particular on the binary mass ratio $q$, disk shape and temperature, and may also depend upon the assumed equation of state. Previous 2D Newtonian simulations, featuring sub-pc fixed binary orbit, predicted $|\xi|\lesssim 2$ (Dittmann & Ryan, 2022) for $H/R\sim 0.03$, although for a higher kinematic viscosity value than the one we simulate in this work.
Note that, due to our live binary setup, we can directly compute the change in SMA and eccentricity from the positions and velocities of the binary components, as well as the accretion rate. We can therefore measure $\xi$ and ${\rm f}_{\rm Edd}$ independently for both gas+PN2.5 and gas+PN2 simulations.
The inclusion of gas in the MBHB system induces perturbations in the evolution of the binary orbital phase. The difference in the number of binary cycles before the merger when it evolves in a gaseous environment is the result of a combination of environmental effects. Indeed the gas causes a different evolution of the binary semi-major axis, eccentricity and precession rate. Since we lack analytical prescriptions for each of these terms in the PN approximation, we can only directly measure the gas-induced orbital ($\delta\phi_{\rm orb}$) and precession ($\delta\phi_{\rm ecc}$) dephasings by comparing the simulations with and without the gaseous disk. We therefore compare the gas+PN2.5 and PN2.5 orbital phases to infer the global effect that the gaseous disk has on the number of binary cycles of a GW-driven MBHB. We then compare the gas+PN2 and the PN2 runs to better isolate the effect of gas on the binary precession rate without any perturbations to the binary induced by GW emission. Indeed in the gas+PN2 and the PN2 setups the binary semi-major axis and eccentricity remain almost constant with time.
Since most of the dephasing is accumulated at larger separations, where the coupling with the disk is stronger, we align the relevant pairs of simulations as much as possible, given the temporal resolution limitation in our simulations, in phase (set to $\pi$) at the final time. The ideal approach would be to align the phases exactly at merger but evolving our systems further would significantly increase the computational cost without adding significant information on the dephasing accumulated at larger separations. We then measure $\delta\phi_{\rm orb}$ by comparing the orbital phases of the appropriate pairs of simulations at the initial time. We measure $\delta\phi_{\rm ecc}$ by comparing the argument of periapsis $\omega=\tan^{-1}(e_{y}/e_{x})$ of the appropriate pairs of simulations, where $e_{x}$ and $e_{y}$ are the $x$ and $y$ components of the binary eccentricity vector $\vec{e}$. Lastly, under the quasi-circular approximation, we can infer the GW dephasing ($\delta\phi_{\rm GW}$) as twice the orbital dephasing, i.e., $\delta\phi_{\rm GW}\approx 2\delta\phi_{\rm orb}^{\rm(GW)}$. We should in principle compute the dephasing in each eccentric GW harmonic, since our binary is eccentric to begin with. However, the scope here is to report a simple estimate of GW dephasing for intuition and the majority of the power would be anyway contained in the fundamental harmonic (i.e at twice the orbital frequency) for our choice of parameters (Peters & Mathews, 1963).
@@ 3 Results
We show the disk morphology in terms of column density at three SMAs – $a=54.5~r_{s}$, $a=52~r_{s}$, and $a=49.5~r_{s}$ – in Fig. 1 for our MR gas+PN2.5 simulation. [^2] The disk morphology for gas+PN2 simulation is similar to the left panel of Fig. 1. The gas morphology is similar, as expected, to the simulations presented in Franchini et al. (2024) with the addition of short-lived mini-disks owing to our higher resolution. We can clearly see the over-density at the cavity edge, i.e. the lump (Shi et al., 2012), precessing around the binary.
@@
We measure the accretion rate in both our simulations and find a mean Eddington ratio of $\bar{{\rm f}_{\rm Edd}}\approx 1.30$ for gas+PN2.5 simulation and $\bar{{\rm f}_{\rm Edd}}\approx 1.50$ for the gas+PN2 case, both consistent with the analytical expectation for the choice of disk mass ($5{\rm M}_{\sun}$) we made and assuming a steady-state $\alpha$-disk with the properties we chose (Frank et al., 2002). The slightly higher accretion rate in the gas+PN2 case is a natural consequence of the fact that the binary does not decouple from the disk because of the lack of GW radiation, therefore the gas can flow inside the cavity and keep feeding the MBHB more effectively.
@@
Figure 1: Column density ($\Sigma$) plots at three SMAs: $54.5~r_{s}$ (left panel), $52~r_{s}$ (middle panel), and $49.5~r_{s}$ (right panel) for the binary evolution under both GW and gas. Here $\Sigma$ varies between $\sim 10^{3}$-$10^{7}$ g/cm2. Both the binary (green dots) and the cavity shrink with time. Moreover, gas inflow inside the cavity creates short-lived mini-disks.
We compute both gravitational and accretion torques directly from the simulations. We find that the magnitude of the accretion torque is $T_{\rm acc}\sim 10^{-2}T_{\rm grav}$. We show the evolution of the parameter $\xi$ in Fig. 2 for gas+PN2.5 simulations as a function of SMA together with the average values of $\xi$ for both gas+PN2.5 ($\bar{\xi}_{\rm{gas+PN2.5}}=-19.5$) and gas+PN2 ($\bar{\xi}_{\rm{gas+PN2}}=-24.4$) cases. Negative torques are consistent with similar aspect ratio HD studies (Tiede et al., 2020; Heath & Nixon, 2020; Dittmann & Ryan, 2022; Franchini et al., 2022). Unsurprisingly, the mean gas effect is stronger when the binary is only evolving due to the gas torques. Since the value of $\xi$ oscillates around its average value over time, one may approximately infer it as a constant parameter within the range of separations explored in this work.